OpenStudy (anonymous):

In Jessica's room, the distance form door D to closet C is 4 feet. The distance from the closet to window W is 6 feet. The distance from the window to the door is 8 feet. On a floor plan of her room, she draws triangle CDW. Order the angles from least to greatest measure.

OpenStudy (anonymous):

sin(DW)/C=sin(CD)/W=sin(CW)/D i think

OpenStudy (anonymous):

you can draw it though

OpenStudy (anonymous):